Solve linear equation:
-3(a+2)+7(-2a-1)=6a-6(3a+5)

Answer by jainpreeja(10)

-3(a+2)+7(-2a-1)=6a-6(3a+5)
-3a-6+-14a-7=6a-18a-30
-3a-14a+18a=-30+13
-17a+12a=-17
-5a=-17
therefore a=-17/-5
a=17/5